The hypotenuse of a right triangle is 75 m long. The length of one leg is 51 m less than the other. Find the length of the legs
Answer by mananth(16946)

since it is right triangle
x^2+(x-51)^2=75^2 ( Pythagoras theorem)
x^2+x^2-102x +2601= 5625
2x^2-102x-3024=0
/2
x^2-51x-1512=0
x^2-72x+21x-1512=0
x(x-72)+21(x-72)=0
(x-72)(x+21)=0
x=72 OR -21
ignore negative
x= 72